I drive 4 miles in 7 minutes @45m/hr,how many ft/s am I driving?
Answer by Theo(13342)

your question states that you drove 4 miles in 7 minutes.
you stated that you were driving at 45 miles per hour.
this would be wrong because 4 miles in 7 minutes equates to an average speed of (4/7) miles per minutes which would equate to an average speed of 60 * (4/7) = 240 / 7) miles per hour which comes out to be 34.28571429 miles per hour.
you did not need to specify the miles per hour so i'll remove that from the problem statement.
you traveled 4 miles in 7 minutes and the question is:
how many feet per second did you travel?
if you travel 4 miles in 7 minutes, then you are traveling 4 * 5280 = 21120 feet in 7 minutes.
since 7 minutes is equal to 7 * 60 = 420 seconds, then you are traveling 21120 feet in 420 seconds.
divide 21120 by 420 and you get an average speed of 50.28571429 feet per second.
this is equivalent to the fraction of (21120/420) feet per second.
to confirm, you know that you traveled 7 minutes which is equal to 7 * 60 = 420 seconds.
you multiply 50.28571429 * 420 to get a distance of 21120 feet and you divide that by 5280 to get a distance of 4 miles.